OpenStudy (anonymous):

how would i find the y-intercept of this equation by hand? p(x)= -(x-3)^2(1-x)^11(2+x)^5

Directrix (directrix):

p(x)= -(x-3)^2 (1-x)^11 (2+x)^5 --------------------------- Let x = 0 and solve for y. all points on the y-axis has x-coordinate zero. y = - (0 - 3)^2 (1-0)^11 (2+0)^5 Crank this out and see what you get. Then, I'll compare with what I got.

OpenStudy (anonymous):

-9(1)(32)=-288 (0,-288) ?

Directrix (directrix):

I got a - 288. y = - (0 - 3)^2 (1-0)^11 (2+0)^5 y =- (-3)^2 * 1 * 2^5 y =-( 9) * 1* 32 y = -288 --> ordered pair (0, -288)